the sum of two numbers is 13. one number is 1 more than twice the other. find the smaller number.
Answer by checkley77(7072)

X+Y=13 OR X=13-Y
X=2Y+1
13-Y=2Y+1
-Y-2Y=1-13
-3Y=-12
Y=-12/-3
Y=4 ANS. FOR THE SMALLER NUMBER.
X+4=13
X=13-4
X=9 ANS. FOR THE LARGER NUMBER.